You will need the data set in the Table below for this question.

Option Earth Moon

Mass/kg   6.0 × 10^24. 7.4 × 10^22


The two bodies are 3.8 × 10^8 m apart


(a) Treating the bodies as point masses, calculate the gravitational field strength at a distance of 3.42 × 108 m from the Earth, on the line that joins them


1
Expert's answer
2021-02-22T10:19:18-0500

"g=G\frac{M}{(r)^2}-G\frac{m}{(R-r)^2}\\
g=6.67\cdot10^{-11}\frac{6\cdot10^{24}}{(3.42\cdot10^{8})^2}\\-6.67\cdot10^{-11}\frac{7.4\cdot10^{22}}{(3.8\cdot10^{8}-3.42\cdot10^{8})^2}\\g=3.4\cdot10^{-6}\frac{m}{s^2}"